4 boys and 2 girls are sitting on the bench
1) what is the number of possible dispositions
2) same question if the boys are on a side and the girls are on the other side
3) same question if each girl is sitting between 2 boys
4) same question if the girls want to sit next to each other

Answer by Edwin McCravy(20081)


4 boys and 2 girls are sitting on the bench
1) what is the number of possible dispositions
6! = 720

2) same question if the boys are on a side and the girls are on the other side
2 ways to choose the general seating arrangement: 

BBBBGG or GGBBBB

4! = 24 ways to arrange the boys
2! = 2 ways to arrange the girls

That's 2*24*2 = 96 ways.  

3) same question if each girl is sitting between 2 boys
3 ways to choose the general arrangement: 

BGBGBB, BGBBGB, or BBGBGB

4! = 24 ways to arrange the boys
2! = 2 ways to arrange the girls

That's 3*24*2 = 144 ways.

4) same question if the girls want to sit next to each other
5 ways to choose the general arrangement: 

GGBBBB, BGGBBB, BBGGBB, BBBGGB, BBBBGG

4! = 24 ways to arrange the boys
2! = 2 ways to arrange the girls

That's 3*24*2 = 144 ways.

Edwin